OpenStudy (anonymous):

Number of pages 50 100 150 200 Cost 1.25 1.85 2.45 3.05 Which equation gives the cost, y, in terms of the number of pages, x? A. y= 0.012x + 49.985 B. y = 0.012x + 0.65 C. y = 83.3x + 0.65 D. y = 0.012x + 1.25

OpenStudy (anonymous):

okay then about this one

OpenStudy (danjs):

notice the increase in x is the same from point to point, and the increase in y is constant, 0.06 more than the last... it is a linear equation y=m*x+b

OpenStudy (danjs):



OpenStudy (danjs):

calculate the slope, and the y-intercept b (initial cost of 0 pages)

OpenStudy (danjs):

the cost is determined by the #pages, so it is independent ly chosen variable, and is usually horizontal in the x direction

OpenStudy (anonymous):

thank you can you help me with this one though

OpenStudy (anonymous):

Which is the equation of the given line in slope-intercept form? a.y= -1/4x+1 b. y= -4x-1 c. y= 4x-1 d. y= 4x+1

OpenStudy (danjs):

slope-intercept form y = m*x+b

OpenStudy (danjs):

use any 2 points on the line to figure the slope m, that one looks lik eyou can read the y-intercept right from the grpah, crosses a gridline on y-axis

OpenStudy (danjs):

what about line equations is the confusing part?

OpenStudy (anonymous):

All of it cause i forgot all about this stuff